### v1

# Promptly Predicting Structures: The Return of Inference

Abstract
Prompt-based methods have been used extensively across NLP to build zero- and few-shot label predictors. Many NLP tasks are naturally structured: that is, their outputs consist of multiple labels which constrain each other. Annotating data for such tasks can be cumbersome. Can the promise of the prompt-based paradigm be extended to such structured outputs? In this paper, we present a framework for constructing zero- and few-shot linguistic structure predictors. Our key insight is that we can use structural constraints—and combinatorial inference derived from them—to filter out inconsistent structures predicted by large language models. We instantiated this framework on two structured prediction tasks, and five datasets. Across all cases, our results show that enforcing consistency not only constructs structurally valid outputs, but also improves performance over the unconstrained variants.
Promptly Predicting Structures: The Return of Inference
1 Introduction
Structured prediction requires making multiple inter-dependent structurally constrained decisions Smith (2010); Nowozin et al. (2014). Many NLP tasks are naturally structured. Consider, for instance, the task of Semantic Role Labeling (SRL) that aims to identify the semantic roles played by sentence constituents for a predicate Palmer et al. (2010). Given a sentence ‘Elrond gave Aragorn the sword.’ and predicate gave, the task entails extracting its semantic arguments — the giver (Agent), the receiver (Recipient), and the thing given (Theme). These decisions are mutually constrained; e.g., the arguments have to be distinct phrases.
Existing methods for structured prediction rely on large labeled datasets. However, obtaining structured labels is not straightforward, especially from non-experts. It invariably requires detailed annotation guidelines about the task, the label set, and the interactions between labels. Methods like the re-formulation of structured prediction tasks into a question-answering (QA) format He et al. (2015); Levy et al. (2017); Du and Cardie (2020) intend to make such tasks more amenable for annotation. However, procuring large-scale annotations still remains expensive and time-intensive; with the problem even more acute for low-resource domains. This predicament necessitates focus on zero and few-shot methods for structured prediction.
Recently, prompt-based methods (e.g., Schick and Schütze, 2021a; Schick and Schütze, 2021b; Liu et al., 2023, inter alia) have shown immense promise, requiring few or even no labeled examples for competitive performance. Prompt-based models generate output text conditionally based on textual input called prompts. Prompts can be crafted using a pre-defined task-specific template Raffel et al. (2020), or even just as natural language instructions Chung et al. (2022) or questions Gardner et al. (2019).
Prompt-based methods have shown promise for text classification and text generation tasks. However, their use for structured prediction is hitherto largely unexplored. Prior work (e.g., see Liu et al., 2023) uses prompts to predict components of structures independently, but their interactions are ignored. Figure 0(a) illustrates a failure of that agenda. Each question pertains to a semantic role for the highlighted predicate. Since the model independently encounters each question, it runs the risk of predicting structurally inconsistent answers.
In this paper, we propose to reintroduce inference into prompt-based methods for structured prediction problems (Figure 0(b)). We present a framework that allows using well-studied inference algorithms with prompt-based predictions. Inference algorithms—realized as beam search, integer linear programs (ILPs, e.g., Roth and Yih, 2004), weighted graph optimization (e.g., Täckström et al., 2015), etc.—take scored candidate label sub-structures as input and optimize a global score while satisfying structural constraints. Historically, trained models generated the label scores (e.g., Lafferty et al., 2001; Collins, 2002). We argue that large language models (LLMs), with appropriate prompts, can be used for scoring, thus eliminating the need for any training or fine-tuning.
We instantiate this idea on two English structured prediction tasks across five datasets. We explore various facets of prompt engineering (e.g., zero- vs few-shot, instruction tuning, etc.) and their impacts on performance. We show that current models can produce structurally inconsistent outputs, and inference helps improve not only consistency, but also overall task performance.111Code available at https://github.com/utahnlp/prompts-for-structures
2 Related Work
Prompts for Predictions.
The “pretrain, prompt, and predict” paradigm (cf. Liu et al., 2023) exploits the observation that LLMs yield general-purpose tools applicable to several natural language tasks. These LLMs (e.g., T5 Raffel et al. (2020), GPT-3 Brown et al. (2020)) may be pretrained for various objectives, including the use of labeled data transformed into prompt-answer pairs. Wei et al. (2022) introduced instruction tuning that transforms natural language tasks into natural language instructions to generalize across tasks independent of their presence in training. With fine-tuning still possible, the prompting paradigm allows for generalization across tasks with some (few-shot) or even no (zero-shot) labeled samples. Several works have shown promise on zero- and few-shot text classification and generation tasks (e.g., Raffel et al., 2020; Brown et al., 2020; Schick and Schütze, 2021b; Lu et al., 2022; Touvron et al., 2023, inter alia).
Structures & Prompts.
Many NLP tasks—e.g., parsing, coreference resolution, information extraction—call for predicting structured outputs such as sequences, trees or graphs. Structured prediction has a rich history in NLP Smith (2010).
Predicting structures with prompts is challenging; the output correspond to multiple mutually-constraining decisions. Structures that can be flattened into label sequences can be cast as text generation tasks amenable to the LLM interface. Blevins et al. (2023) present a structured prompting paradigm to generate labels in an auto-regressive fashion for part-of-speech tagging, named entity recognition, and sentence chunking tasks. They show that this approach works well in a few-shot setting. Such tasks are also amenable for templated slot filling in a few shot setting Cui et al. (2021). Liu et al. (2022) propose an autoregressive structured prediction framework that reformulates the problem into predicting a series of structure-building action steps.
Structured prediction tasks can be made prompt-friendly by identifying output components and framing each one into natural instructions He et al. (2015); Du and Cardie (2020) using the QA format. Mekala et al. (2023) decompose semantic parsing into a series of questions. However, they do not use inference and instead, aggregate outputs sequentially. Yang et al. (2022) use QA prompts to classify pairs of entity mentions as coreferent or not without aiming to form entity clusters. The near contemporary work of Lin et al. (2023) presents a prompt-driven heuristic for event type classification. Wang et al. (2023) use a programming language-based prompt and generate approach with GPT-3 for zero and few-shot event argument extraction. Recently, Le and Ritter (2023) show two techniques based on QA and document infilling for zero-shot coreference resolution that rely on recent LLMs’ ability to ingest entire documents as input.
3 “Promptly” Predicting Structures
Problem Statement and Notation
Given an input (e.g., phrase, sentence, etc.), the goal of a conditional model is to characterize the distribution over outputs . The outputs in structured prediction tasks consist of a set of decisions . For instance, consider the SRL task from §1. Given a predicate, identifying the token spans correponding to its arguments requires multiple decisions; i.e., each corresponds to a semantic argument.
Since the size of the output is variable and depends on the input , the standard approach for modeling such problems calls for factorizing the distribution over its components . The factorization is a design choice; the simplest one decomposes the probability as a product over individual local decisions: . With such a factorization, the prediction for an input is given by
These local probabilities relate to unary potentials in factor graph notation, and in our work, will correspond to logits from neural models. Importantly, the maximization problem above is trivial to solve. Each can be independently assigned.222Of course, the structured prediction literature studies other task-dependent factorizations. With neural models, auto-regressive factorizations are also common, where each label is conditioned on previously predicted ones. In this work, we study the simple factorization described here because it can be easily adapted to work with prompting.
Even the simple factorization above presents two drawbacks: i) It requires an explicit training step using expensive-to-acquire training data to obtain the unary potentials. ii) It does not guarantee structural consistency, leading to invalid structures. In this work, we ask: Can we predict valid structured outputs in the zero- to few-shot regime?
Now, let us address these drawbacks, leading to our proposed framework.
Unary Potentials from Prompts.
Prompt-based methods can help bypass the need for supervised training. For every local decision , we construct a query to convert the decision into a prompt-friendly format. A pretrained model can use these queries in the standard prompt-and-predict fashion to get scored candidate answers. Formally, we write . Here, is a set of questions corresponding to the components of the structure. In the case of SRL, would be the input sentence and the predicate, and would contain one question per semantic argument type for that predicate. Figure 1 gives an example.
The Return of Inference.
Merely aggregating outputs generated with prompts does not address the second drawback, i.e., the need for structurally valid outputs. Structural validity is a constraint that, when applied to the prediction problem, ensures that only valid output collections are considered as candidate predictions. In other words, instead of the unconstrained prediction of Eq. 1, we have
The inference problem above requires combinatorial search over valid structures. Inference can be realized algorithmically in various forms like beam search, an ILP Roth and Yih (2004) or weighted graph search (e.g., Täckström et al., 2015; Zaratiana et al., 2022).
The Framework.
We propose the following blueprint for predicting valid structures in a zero-shot manner: i) Break the prediction task into its constituent sub-tasks. ii) Convert each sub-task into a prompt-friendly format. iii) Implement an inference algorithm that takes scores from the LLM and produces a valid structured output. This recipe naturally extends to a few-shot setting by prefixing in-context examples with the original prompt.
Inference constructs outputs that satisfy constraints inherent in the definition of a task. Moreover, it can help correct errors of the zero- or few-shot models using other predictions of the same model. Without this mechanism (i.e., using the labels from the prompts directly as in Equation 1), not only will we have partially wrong outputs, they may be invalid from the perspective of the task. Finally, as we will see in the case of the coreference task, inference can help make long-range decisions about parts of an input even when the entire input is not presented to an LLM. In this sense, it can be seen as a mechanism to “bring back the context” without explicitly having to use a long-document model such as LongFormer Beltagy et al. (2020).
4 Experiments
This section describes the tasks and datasets where we instantiate the framework from section §3.
4.1 Semantic Role Labeling
Semantic role labeling (SRL) is the task of constructing predicate-argument structures triggered by verbs in a sentence. This task has been studied in using the PropBank Palmer et al. (2005), FrameNet Baker et al. (1998), and more recently, the QA-SRL He et al. (2015) datasets.
Datasets.
The QA-SRL project He et al. (2015) recasts the argument labeling task as question-answering to make it amenable for non-expert annotation. By breaking the predicate-argument structure into a collection of individual arguments, the project curated a collection of QA pairs that represent SRL frames. This SRL-to-QA transformation sets it up perfectly for prompting; the sentence along with each question can serve as a prompt.
We consider two QA-SRL datasets for our experiments. 1) QA-SRLHe et al. (2015)contains 10.8k QA pairs for 4440 predicates in 1959 sentences across splits.333He et al. (2015) provide datasets in newswire and Wikipedia domains. We use the latter as it is larger. 2) FitzGerald et al. (2018)present a much larger crowd-sourced dataset—QA-SRL 2.0—containing 265k QA pairs from 64k sentences for 133k verbal predicates annotated using the QA-SRL annotation schema.
Constraints.
The SRL definition dictates that different argument spans for a predicate do not have token overlap Palmer et al. (2010), and the QA-SRL schema mandates an answer per question.
Prompts and Inference.
For any predicate in a sentence, we have one question per semantic role . For every question, we prompt a generative language model to predict the best- spans (i.e., sequences) with their respective scores. We denote the score from the language model for the (where r n) ranked span () for role as . The scored spans from all questions for a predicate are input to an inference algorithm.
We follow the inference algorithm of Täckström et al. (2015).444For brevity, we outline the inference setup here, and refer the reader to the original work for a detailed description. Given a sentence, we construct a directed graph with one more vertex than the number of tokens. (Figure 2 shows an example.) A vertex is shared by two consecutive words and , and denotes the boundary between them. Edges can be of two kinds: i) Null edges are added between consecutive nodes with a zero edge score. ii) For every candidate span ranked for the role i, we add an edge between its representative vertices with an edge score . The highest ranked span gets an edge score of zero.555A generative model tasked for extraction, like ours, cannot get indices to mark the start and end. Yet, a generated span may occur multiple times in a sentence. In such cases, edges representing all such spans are added with an identical score.
Given the graph, we can perform a shortest-path search from the leftmost to the rightmost nodes; the edges in the shortest path gives us non-overlapping labels. However, doing so will generate null sequences that have a zero score. We avoid this issue by extending the algorithm to return the top shortest paths using Yen’s -shortest path algorithm Yen (1971). Subsequently, the highest scoring path among the paths which assigns exactly one span per semantic role is the optimal structure that satisfies all task constraints. For the example in Figure 2, our inference algorithm would return ‘Elrond’, ‘Aragorn’, and ‘the sword’ as the arguments for roles a, b and c respectively.666See Appendix F for a worked-out example.
Evaluation.
We evaluate at both the question-level and the structure-level. At each level, we use exact and head match metrics (Palmer et al., 2010, p. 49). At the question-level, predicted spans are compared with gold spans. For the Exact accuracy metric, answers are considered correct if the spans exactly match. Similarly, answers are considered correct if the phrasal heads of the spans match to compute the Head accuracy metric. A structure is considered accurate if predicted spans for all constituent roles match. As at the question-level, a match can be determined in an exact (Exact), or head (Head) fashion. Additionally, we also evaluate inconsistency percent () by measuring how often constraints are violated. We compare every pair of outputs(i.e., argument spans) in a predicted structure. A violation occurs when two output spans overlap.
4.2 Coreference Resolution
Coreference resolution is the task of grouping mention expressions (or simply, mentions) in a document into clusters that refer to the same entity. For example, consider the sentence “Al requested Bob to give him the pen.”, with given mentions: “Al”, “Bob”, “him”, “the pen”. Here, “Al” and “him” refer to the same entity, i.e., Al, while the other mentions are singletons.
Datasets.
We use three datasets for this task: 1) The entity coreference annotations within the EventCorefBank+ dataset (ECB+, Cybulska and Vossen, 2014).777We use the processed files provided by Yang et al. (2022) 2) The CoNLL 2012 OntoNotes 5.0 dataset with the v12 English data splits Pradhan et al. (2012), which contains entity and event coreference annotation across several domains like magazine articles, broadcast conversations, among others. 3) The GENIA Coreference Corpus Su et al. (2008), which helps showcase the applicability of this framework in a low-resource domain like biomedicine where annotation is especially difficult. The GENIA corpus is a collection of coreference annotations for 2000 paper abstracts published in the biomedical domain. We define our own train/dev/test splits for the corpus since pre-defined splits were unavailable. We will release these splits for future use. For OntoNotes and GENIA, we do not consider nested entities.
Constraints.
The coreference resolution task can be broken into a series of binary decisions: for every mention pair in a document we can ask whether the mentions are co-referrant or not. Entity clusters require transitive closure. That is, given three mentions , and , if a predicate denotes and refer to the same entity, then:
Prompt and Inference.
For every document, we collect mention pairs in it.888For Ontonotes, we considered all mention pairs in a window size of three sentences due to a large number of mentions. We transform the coreference decision about mentions and into a Yes/No question: “Does refer to ?”. A ‘Yes’ establishes a coreference link between the mentions. Questions are appended with the relevant context and given to a language model. Link scores () for a mention-pair and are given by where and are the scores for generating the sequences "yes" and "no" given a question and context. Using the link scores, we employ All-link inference from Chang et al. (2011) to solve the following integer program over binary decisions :
| such that | |||
The constraint ensures that the transitivity condition from (3) is satisfied.
Evaluation.
We use two metrics: i) F1that measures performance of the binary decisions at the question level ii) CoNLL scorePradhan et al. (2014)that averages three cluster evaluation metrics: MUC , B, and CEAF. Additionally, we measure the inconsistency percent (): the number of times constraint (3) is violated divided by the number of times the antecedent is true. This is the same as the ‘conditional violation’ defined by Li et al. (2019), and measures the fraction of times a third link between three mentions does not exist, when the other two links between them exist.
5 Results
This section presents the impact of inference on zero-shot performance and consistency. We use the 3 billion variants from the T5 Raffel et al. (2020) family of encoder-decoder models for our experiments. We discuss our implementation briefly here; Appendix D has details like prompt templates.
5.1 Semantic Role Labeling
We consider the T5-3B model999We also considered UnifiedQAv2 Khashabi et al. (2022) and Macaw Tafjord and Clark (2021). T5 performed the best. for our experiments. We also present results on Flan T5-XL Wei et al. (2022), also a 3B parameter model, to check if instruction tuning helps with zero-shot performance.
Instruction-tuned models like Flan T5 allow verbalizing constraints within the prompt structure. We also test out an iterative prompting method Wang et al. (2022) where questions pertaining to a predicate in a sentence are presented in sequential order to the model with verbalized constraints. In every prompt, we include the QA pair from every previous question for the predicate already prompted. This prompting strategy is denoted by the superscript itr. See Appendix D.1.1 for an example.101010Note that verbalized constraints do not guarantee a valid output structure. However, exposure to preceding predictions may help reduce inconsistencies nonetheless. Further, we can use inference in conjunction with this iterative prompting strategy.
Tables 1 and 2 show the results for QA-SRL and QA-SRL 2.0 respectively. We see that the constrained models produce near-consistent gains over their unconstrained counterparts. Flan-T5 does not outperform the T5 model, but iterative prompting with inference significantly improves its performance. More importantly, without inference, the models produce highly inconsistent predictions with a lowest inconsistency percent of 34%! In addition, we also provide results on GPT-4 with the constraints verbalized as system instruction.111111GPT-4 results are meant to contextualize our results with a current SOTA system. The datasets we considered may have been utilized to train GPT-4. GPT-4 also has multiple orders of magnitude more parameters than the models we consider. For GPT-4, all the questions are provided at once. Refer Appendix D.1.2 for details.
As an illustrative example, we present a case where constrained inference corrects overlapping predicted arguments. For the sentence, “Keats’ long and expensive medical training with Hammond and at Guy’s Hospital led his family to assume he would pursue a lifelong career in medicine, assuring financial security …”, the model was asked: i) What would assure something? ii) What would something assure? The unconstrained model generated “financial security” on both counts. Post-inference, the answer to the first question is changed to “lifelong career in medicine”, resulting in the correct structure.
| System | Head() | Head() | () |
| T5-3B | 63.84 | 36.17 | 34.30 |
| + constraints | 62.11 | 38.75 | 0 |
| Flan-T5-XL | 36.76 | 11.20 | 87.16 |
| + constraints | 43.34 | 18.48 | 0 |
| Flan-T5-XL | 59.47 | 33.60 | 49.79 |
| + constraints | 61.11 | 36.17 | 0 |
| GPT-4 | 80.46 | 62.93 | 5.61 |
| System | Head() | Head() | () |
| T5-3B | 68.57 | 46.35 | 39.06 |
| + constraints | 68.87 | 52.31 | 0 |
| Flan-T5-XL | 61.92 | 40.37 | 58.19 |
| + constraints | 66.02 | 48.68 | 0 |
| Flan-T5-XL | 68.84 | 49.29 | 42.11 |
| + constraints | 71.97 | 55.41 | 0 |
| GPT-4 | 86.78 | 76.89 | 7.18 |
5.2 Coreference Resolution
The ability to specify multiple choices makes the Macaw-3B Tafjord and Clark (2021) model an ideal choice for the Yes/No choice QA task.121212We also considered this task in an NLI format with the T5-3B model. However, the QA format was a clear winner. As before, we show results using the instruction-tuned Flan T5-XL model as well. For this set of experiments, we only prompt the model with the sentence(s) containing the mention pair as context, and not the whole document.131313Note: We do not need a large context window. Inference can create consistent clusters by looking at smaller contexts. We show more analysis with varying context styles in Appendix B.2.
| System | F1() | CoNLL() | () |
| All-Yes | 13.32 | 41.51 | 0 |
| All-No | 45.84 | 45.06 | 0 |
| Macaw-3B | 46.26 | N/A | 48.56 |
| + R2L | 42.20 | 53.55 | 0 |
| + All-Link | 48.57 | 57.76 | 0 |
| Flan-T5-XL | 61.58 | N/A | 80.78 |
| + R2L | 58.38 | 62.79 | 0 |
| + All-Link | 66.06 | 65.09 | 0 |
In addition to the constrained inference proposed in §4.2, we consider a simple right-to-left cluster assignment heuristic (R2L) proposed by Soon et al. (2001). This heuristic assigns a candidate mention to the cluster of the closest mention where a link was predicted by the unconstrained model. We also present two rudimentary baselines: i) All-Yesbaseline where all questions are answered “Yes” and all mentions belong to a single cluster, and ii) All-Nobaseline where all questions are answered “No” and every mention is its own cluster. These baselines are relevant since both generate valid structures.
The unconstrained and the constrained models can only be compared using their F1 scores and the inconsistencies.141414The official CoNLL scorer requires valid structures rendering base system evaluation impossible w/o inference. Tables 3, 4 and 5 show results for the ECB+, OntoNotes and GENIA datasets respectively. We refer the reader to Table 8 in Appendix B.1 for a detailed breakdown of the CoNLL score. We observe gains in F1 scores for constrained systems over their unconstrained counterparts. In all cases, our models outperform the All-Yes and All-No baselines. We also see that the inconsistency is substantially higher for the unconstrained models, especially with Flan-T5-XL.
| System | F1() | CoNLL() | () |
| All-Yes | 18.96 | 38.61 | 0 |
| All-No | 43.38 | 39.42 | 0 |
| Macaw-3B | 48.67 | N/A | 63.88 |
| + R2L | 49.48 | 55.30 | 0 |
| + All-Link | 52.15 | 55.14 | 0 |
| Flan-T5-XL | 48.93 | N/A | 86.75 |
| + R2L | 51.88 | 50.37 | 0 |
| + All-Link | 51.33 | 48.52 | 0 |
| System | F1() | CoNLL() | () |
| All-Yes | 9.00 | 37.70 | 0 |
| All-No | 47.40 | 31.19 | 0 |
| Macaw-3B | 46.63 | N/A | 51.18 |
| + R2L | 35.62 | 47.25 | 0 |
| + All-Link | 46.75 | 52.56 | 0 |
| Flan-T5-XL | 56.53 | N/A | 82.64 |
| + R2L | 53.35 | 54.11 | 0 |
| + All-Link | 65.36 | 57.47 | 0 |
Figure 3 presents an instance where inference corrects flawed links. Here, ‘T-mobile’ and ‘carrier’ belong to one entity cluster, while, ‘Blackberry Curve 8900’ and ‘product’ belong to another. Our unconstrained model incorrectly predicts a link between ‘T-mobile’ and ‘Blackberry Curve 8900’. But transitivity requires all other edges to be present, which the model does not predict. Our inference algorithm removes the erroneous edge and predicts consistent clusters.
Iterative prompting for coreference resolution is challenging for two reasons: 1) As the component questions establish links between two mentions and the transitivity closure applies to three mentions, verbalizing the transitivity constraint within the prompt can be challenging. 2) Coreference structures are much larger than the SRL structures, hence, demanding a large memory requirement.
6 Analysis Experiments
Next, we show how performance and inconsistency vary with model size and in-context examples.
6.1 Model Size
We study the performance and consistency of models over varying sizes. For SRL, we compare the small, base, large, 3 billion, and 11 billion T5 models. Figure 4 shows the Head metric over T5 model sizes for QA-SRL 2.0, along with the inconsistency percent. For coreference, we compare the large, 3 billion and 11 billion variants of the Macaw model. We consider the All-Link inference for the constrained models. We report the F1 score and the inconsistency percent on the OntoNotes dataset in Figure 5. The inconsistency is shown only for the unconstrained models; the constrained ones are guaranteed to be consistent.
Across the datasets, we see that performance almost always increases and constraints provide steady gains. In QA-SRL 2.0, inconsistency reduces with increasing model size. The trend reverses for OntoNotes. Interestingly, with QA-SRL 2.0, a constrained model of a certain size often yields comparable, if not better, results as an unconstrained model of the immediately bigger model, showing that constraints might give performance gains equivalent to using a bigger model. Appendix C shows model size results for the other datasets.
6.2 Few-shot Prompting
So far, we have examined our framework in the zero-shot setting. Do the conclusions hold for few-shot settings as well? To answer this question, we use the Flan-T5-XL model which is known to yield promising results with even a few in-context examples.151515An interesting question: What counts as a shot? All questions in a structure or just one? We consider the latter. We conduct few-shot experiments for the SRL task on the QA-SRL (Figure 6) and QA-SRL 2.0 (Figure 7). We observe a clear increase in performance with increasing shots for QA-SRL. Recall that Flan T5, in a zero-shot setting for this dataset, underperforms compared to T5. Hence, the gains are not surprising: only the 5-shot constrained system beats the zero-shot constrained T5 model (38.75 Head). The model also becomes more consistent with more examples. For QA-SRL 2.0, the performance and the consistency largely remain invariant with number of shots. Constrained systems produce more accurate structures than their unconstrained counterparts.
7 Conclusions
Prompt-based methods have impressed when deployed in zero- and few-shot manner. Structured prediction problems can use them by breaking the structure down into smaller components and querying for local decisions. This aspect has been explored by some in the literature. However, these works leave out a pivotal part of such structured prediction tasks—the structure. In this work, we proposed a new framework to predict structurally consistent outputs by following local predictions with an inference step. In our experiments, our framework not only guarantees consistency but also provides consistent performance gains.
Limitations
Our focus of this work is to predict structures given that a structured prediciton task can be broken down into components and questions can be generated for those components. While this relies heavily on advances in question generation, we do not focus on this line of research and consider that questions are given to us. A separate line of work exists precisely in this domain of question generation for structured prediction tasks He et al. (2015); Levy et al. (2017); Du and Cardie (2020); Wu et al. (2020). We emphasize that our contribution is more general as to how we can marry the idea of inference with prompting.
A problem that can affect any generative model is the label bias problem where the model prefers a certain label over certain other label irrespective of the inputs. This problem can cause performance imbalances and models need to be calibrated to alleviate this phenomenon. Recently, a few works Zhao et al. (2021); Holtzman et al. (2021); Chen et al. (2023); Han et al. (2023); Fei et al. (2023) study this phenomenon for zero and few shot prompting models for an array of text classification tasks. We briefly studied this for our coreference experiments and found that Macaw was more well-calibrated for the binary task as compared to T5. We leave a detailed calibration study for future work.
Acknowledgements
We thank the members at the Utah NLP lab and AI2 for their insights, especially, Ashim Gupta for his feedback on the iterative prompting experiments. We thank the anonymous reviewers for their feedback. This material is based upon work supported in part by NSF under grants #2007398, #2217154, #1822877 and #1801446. The support and resources from the Center for High Performance Computing at the University of Utah are gratefully acknowledged.
Appendix A Detailed SRL Results
| System | Exact() | Exact() | Head() | Head() | () |
| T5-3B | 38.80 | 12.65 | 63.84 | 36.17 | 34.30 |
| + constraints | 40.30 | 14.78 | 62.11 | 38.75 | 0 |
| Flan-T5-XL | 22.72 | 3.47 | 36.76 | 11.20 | 87.16 |
| + constraints | 33.12 | 9.52 | 43.34 | 18.48 | 0 |
| Flan-T5-XL | 37.57 | 12.32 | 59.47 | 33.60 | 49.79 |
| + constraints | 43.80 | 17.25 | 61.11 | 36.17 | 0 |
| GPT-4 | 62.84 | 37.07 | 80.46 | 62.93 | 5.61 |
| System | Exact() | Exact() | Head() | Head() | () |
| T5-3B | 54.12 | 30.49 | 68.57 | 46.35 | 39.06 |
| + constraints | 56.78 | 36.96 | 68.87 | 52.31 | 0 |
| Flan-T5-XL | 48.92 | 26.53 | 61.92 | 40.37 | 58.19 |
| + constraints | 57.06 | 37.23 | 66.02 | 48.68 | 0 |
| Flan-T5-XL | 55.11 | 32.86 | 68.84 | 49.29 | 42.11 |
| + constraints | 61.34 | 41.17 | 71.97 | 55.41 | 0 |
| GPT-4 | 77.98 | 63.00 | 86.78 | 76.89 | 7.18 |
Appendix B Additional Coreference Results
B.1 Zero Shot Coreference Results
The complete zeros shots results for Macaw-3B and Flan T5-XL models are given in Table 8. One might notice that the Flan model is highly inconsistent as compared to the Macaw model. Note that in each case, the divisor is different since it is the number of times the antecedent is activated in equation 3. Examining the predictions, we saw that the value of the divisor for the Macaw models is roughly four times that of the Flan models. This implies that Flan tends to predict ‘No’ more often.
| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
| Single | 13.32 | 63.93 | 41.67 | 18.94 | 41.51 | 0 | |
[…]
| ECB+ | + R2L | 42.20 | 52.71 | 61.00 | 46.93 | 53.55 | 0 |
| + All-Link | 48.57 | 56.72 | 65.59 | 50.98 | 57.76 | 0 | |
| Flan-T5-XL | 61.58 | N/A | N/A | N/A | N/A | 80.78 | |
| + R2L | 58.38 | 53.03 | 73.27 | 62.08 | 62.79 | 0 | |
| + All-Link | 66.06 | 51.59 | 77.07 | 66.60 | 65.09 | 0 | |
| Single | 18.96 | 70.23 | 30.89 | 14.71 | 38.61 | 0 | |
[…]
| Ontonotes | + R2L | 49.48 | 51.11 | 66.04 | 48.76 | 55.30 | 0 |
| + All-Link | 52.15 | 47.67 | 67.02 | 50.72 | 55.14 | 0 | |
| Flan-T5-XL | 48.93 | N/A | N/A | N/A | N/A | 86.75 | |
| + R2L | 51.88 | 31.12 | 67.51 | 52.47 | 50.37 | 0 | |
| + All-Link | 51.33 | 24.93 | 67.08 | 53.56 | 48.52 | 0 | |
| Single | 9.00 | 77.41 | 26.81 | 8.89 | 37.70 | 0 | |
[…]
| GENIA | + R2L | 35.62 | 66.00 | 45.11 | 30.65 | 47.25 | 0 |
| + All-Link | 46.75 | 62.40 | 56.19 | 39.08 | 52.56 | 0 | |
| Flan-T5-XL | 56.53 | N/A | N/A | N/A | N/A | 82.64 | |
| + R2L | 53.35 | 59.62 | 58.30 | 44.41 | 54.11 | 0 | |
| + All-Link | 65.36 | 53.48 | 66.90 | 52.04 | 57.47 | 0 |
B.2 Context Styles for Coreference Resolution
As mentioned in §5.2, for any given question pertaining to a mention-pair, we consider only the sentences where the constituent mentions occur. One can argue that surrounding context can play an important role in deciding whether or not two mentions are co-referent. Furthermore, intermediate context between the sentences can be pivotal in establishing otherwise loosely connected sentences appearing a few sentences apart. This begs the question about how much, if any, does providing the entire document help.161616For Ontonotes and GENIA, the document size went beyond the input size which is why we do not test along this dimension. On a different note, mentions can be highlighted with asterisks(as used for WSC in Raffel et al. (2020)) to explicitly state where in the context they occur. This approach can help deal with pronominal mentions which are bound to repeat. We present results when the entire document (Full) acts as context, when mentions are highlighted (Hlght), and a combination of both (Full + Hlght) in comparison to the context style where only the relevant sentences are considered (Rel). We provide detailed results for performance and consistency over differing context styles in Table 9. We see that results are comparable across all context styles.
| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
| Rel | 46.26 | N/A | N/A | N/A | N/A | 48.56 | |
| ECB+ | + R2L | 42.20 | 52.71 | 61.00 | 46.93 | 53.55 | 0 |
| + All-Link | 48.57 | 56.72 | 65.59 | 50.98 | 57.76 | 0 | |
| Hlght | 46.03 | N/A | N/A | N/A | N/A | 46.68 | |
| + R2L | 42.49 | 51.19 | 60.76 | 47.17 | 53.04 | 0 | |
| + All-Link | 47.84 | 57.92 | 64.80 | 49.59 | 57.44 | 0 | |
| Full | 44.47 | N/A | N/A | N/A | N/A | 31.15 | |
| + R2L | 38.46 | 57.51 | 60.25 | 43.97 | 53.91 | 0 | |
| + All-Link | 45.59 | 59.28 | 64.69 | 48.93 | 57.63 | 0 | |
| Full+Hlght | 43.54 | N/A | N/A | N/A | N/A | 31.02 | |
| + R2L | 37.38 | 59.17 | 59.44 | 43.15 | 53.92 | 0 | |
| + All-Link | 44.66 | 60.34 | 63.80 | 48.56 | 57.57 | 0 | |
| Rel | 48.67 | N/A | N/A | N/A | N/A | 63.88 | |
| Ontonotes | + R2L | 49.48 | 51.11 | 66.04 | 48.76 | 55.30 | 0 |
| + All-Link | 52.15 | 47.67 | 67.02 | 50.72 | 55.14 | 0 | |
| Hlght | 49.22 | N/A | N/A | N/A | N/A | 64.93 | |
| + R2L | 49.40 | 51.16 | 65.88 | 48.75 | 55.26 | 0 | |
| + All-Link | 52.43 | 48.20 | 67.13 | 50.90 | 55.41 | 0 | |
| Rel | 46.63 | N/A | N/A | N/A | N/A | 51.18 | |
| GENIA | + R2L | 35.62 | 66.00 | 45.11 | 30.65 | 47.25 | 0 |
| + All-Link | 46.75 | 62.40 | 56.19 | 39.08 | 52.56 | 0 | |
| Hlght | 46.55 | N/A | N/A | N/A | N/A | 50.62 | |
| + R2L | 34.94 | 66.29 | 45.30 | 30.81 | 47.47 | 0 | |
| + All-Link | 46.66 | 62.67 | 55.92 | 39.22 | 52.60 | 0 |
Appendix C Additional Model Size Experiments
The analysis over model sizes for QA-SRL is given in Figure 8. We see similar analysis as observed with QA-SRL 2.0 dataset. Performance improves with increasing model size while inconsistency reduces. Furthermore, constraints provide steady gains over their unconstrained counterparts.
Figure 9 shows performance and inconsistency of the Macaw model over sizes for the ECB+ coreference dataset. Figure 10 shows the same for the GENIA dataset. In both the cases, we see that performance drastically improves from the large to the 3 billion parameter variant. Performance gains from 3 billion to 11 billion are minimal. However, we see that the inconsistency increases with increasing size. We also see that for these datasets, for the large variant, constraints do not help the model. However, note that F1 is a question-level metric and a drop might be a necessary to guarantee a consistent structure.
Appendix D Experimental and Prompt Details
We use HuggingFace’s transformers library Wolf et al. (2020) for all our models. All model generations were done with a random seed of 2121. We use the ‘generate’ method in the HuggingFace Wolf et al. (2020) library to return the score of a generated sequence. These scores can be accessed by the ‘sequences_scores’ output attribute. We consider the top 20 spans for each question. For all few shot settings, we average results across shots considered from three seeds: 42, 20 and 1984.
We explain the prompt templates and the parameters/assumptions we chose in the subsequent sections.
D.1 Semantic Role Labeling.
[…]
D.1.1 Iterative Prompting Template
In the iterative prompting template, we prompt the model one question at a time such that the answer for every preceding prompt is included in subsequent prompts. Assume a predicate in a sentence to have questions: {q,q,…,q}. In the first step, we prompt the model for q using the prompt template mentioned in the previous sub-section. Next, we follow with the prompt the model for q such that the question and the generated answer for q are included in the prompt. Consequently, we use a different prompt template that contains this pair and an instruction that asks the answers to not overlap. Subsequently, the prompt for q will contain the question-answer pair q and q, and so on. An example is shown in Figure 11.
D.1.2 GPT-4 Prompt Template
We used OpenAI’s API and the ChatCompletion end-point to obtain the GPT-4 results. We provide the following system instruction: “You will be given a set of questions regarding a particular sentence. Answer these questions such that the answers to the questions do not overlap. Answers should strictly be a sub-sequence of the sentence. Do not include anything except the answer phrase in the answer”. A user prompt consists of a sentence followed by a set of questions pertaining to a certain predicate in the sentence. As a result, each predicate consists of a request to the endpoint. An example request and answer are shown in Figure 12. A majority of the GPT-4 generations adhere to the answer template shown in the figure. However, for predicates with a single question, answers did not always contain answer headers ( e.g., “Answer 1: <answer>”). These were handled accordingly. In some rare cases, “None of the above choices” was generated and such cases were considered as having no answers. The experiment across the two datasets cost a total of $96.31.
D.2 Coreference Resolution.
We are given the context (<context>) and a pair of mentions(<m1> and <m2>) for the input. For the Macaw models, the input prompt follows this format:
“$answer$ ; $mcoptions$=(A) Yes (B) No ; <context> Does <m1> refer to <m2>?”
For the Flan models, we follow the format:
[…]
We restrict generations to {“Yes”,“No”} for Flan experiments and to {“$answer$ = Yes”, “$answer$ = No”} for the Macaw experiments. Integer Linear Programs were implemented and solved using the Gurobi solver Gurobi Optimization, LLC (2023).
Appendix E Inference Overhead
Since the inference algorithm follows the generation step, a time overhead is expected. Table 10 shows the generation time and its corresponding inference time for the datasets we consider. Crucially, while the generation step is run on the GPU, our inference implementation is run on the CPU.
| Dataset | Gen. step | Inf. Step |
| QA-SRL | 48.49 | 2.18 |
| QA-SRL 2.0 | 945.22 | 32.68 |
| ECB+ | 81.87 | 0.08 |
| Ontonotes | 1547.03 | 62.99 |
| GENIA | 859.88 | 116.33 |
Appendix F SRL Inference Worked-out Example
We will consider the example mentioned in Figure 2 with toy values to illustrate the inference process. In this toy setup, we consider three roles (a,b, and c)171717You may assume these to be the Agent, Recipient and Theme roles as mentioned in §1 and assume the top-2 candidate spans per role with their scores as returned a hypothetical generative model.
The spans and scores for this toy example are given in Table 12. Given these spans and scores. we can construct the directed graph as shown in Figure 13. We calculate the edge scores where is the score (as returned by the scoring function) for the top ranked span for role . For the sake of this illustration, we shall use as the span notation as well. The value is set to zero. The first step of our inference algorithm aims to find the shortest paths from the leftmost to the rightmost node. Observe how overlapping spans can never be on the same path. As a results, the top ranked candidates for roles (span ) and (span) can never be in the same structure.
shortest paths are returned in increasing order of total path length where path length for a path S is as shown in Table 11. As a result, one might notice that partial structures such as paths or might have the shortest path length. However, the QA-SRL framework necessitates one answer per question(or, in other words, role). A post-processing step follows the shortest path step where, from the shortest paths returned, the shortest path containing exactly one span for each role is chosen. This step also ensures that paths where multiple spans for the same role are present (e.g., ), are also avoided. In the example shown in Figure 13, gives the most optimal path with a path length 1 and satisfies all the constraints. This yields to “Elrond”, “Aragorn” and “the sword” as arguments for roles a, b and c respectively. Intuitively, a sufficiently large (in our case, 20) leads to a higher chance of obtaining a complete and consistent structure. In cases where no complete structures can be formed from any of the shortest paths, partial (yet, consistent) structures are predicted.181818Partial structures can be considered as having ”null” answers for some roles. For evaluation purposes, we treat them as incorrect answers.
| Rank | Path | Path length |
| 1 | 0 | |
| 2 | 0 | |
| 3 | 0 | |
| 4 | 0 | |
| 5 | 0 | |
| 6 | 0 | |
| 7 | 1 | |
| 8 | 1 | |
| Rank | Role | ||
| a | b | c | |
| 1 | “Elrond” (2) | “Aragorn” (5) | “Aragorn the sword” (5) |
| 2 | “Elrond gave” (1) | “Elrond” (3) | “the sword” (4) |

### v2

@@
 Structured prediction has a rich history in NLP Smith (2010).
-Predicting structures with prompts is challenging; the output correspond to multiple mutually-constraining decisions.
+Predicting structures with prompts is challenging; the output corresponds to multiple mutually-constraining decisions.
 Structures that can be flattened into label sequences can be cast as text generation tasks amenable to the LLM interface.
@@
 Recently, Le and Ritter (2023) show two techniques based on QA and document infilling for zero-shot coreference resolution that rely on recent LLMs’ ability to ingest entire documents as input.
+The concurrent work of Rajaby Faghihi and Kordjamshidi (2024) introduces an inference-time method for enforcing structural consistency that incorporates local decisions from different models using various normalizer functions.
 3 “Promptly” Predicting Structures
@@
 We consider two QA-SRL datasets for our experiments.
-1) QA-SRLHe et al.
+1) QA-SRLwikiHe et al.
 (2015)contains 10.8k QA pairs for 4440 predicates in 1959 sentences across splits.333He et al.
@@
 In such cases, edges representing all such spans are added with an identical score.
-Given the graph, we can perform a shortest-path search from the leftmost to the rightmost nodes; the edges in the shortest path gives us non-overlapping labels.
+Given the graph, we can perform a shortest-path search from the leftmost to the rightmost nodes; the edges in the shortest path give us non-overlapping labels.
 However, doing so will generate null sequences that have a zero score.
@@
 At the question-level, predicted spans are compared with gold spans.
-For the Exact accuracy metric, answers are considered correct if the spans exactly match.
-Similarly, answers are considered correct if the phrasal heads of the spans match to compute the Head accuracy metric.
+For the Exactq accuracy metric, answers are considered correct if the spans exactly match.
+Similarly, answers are considered correct if the phrasal heads of the spans match to compute the Headq accuracy metric.
 A structure is considered accurate if predicted spans for all constituent roles match.
-As at the question-level, a match can be determined in an exact (Exact), or head (Head) fashion.
+As at the question-level, a match can be determined in an exact (Exacts), or head (Heads) fashion.
 Additionally, we also evaluate inconsistency percent () by measuring how often constraints are violated.
@@
 We use two metrics: i) F1that measures performance of the binary decisions at the question level ii) CoNLL scorePradhan et al.
-(2014)that averages three cluster evaluation metrics: MUC , B, and CEAF.
+(2014)that averages three cluster evaluation metrics: MUC , B3, and CEAFe.
 Additionally, we measure the inconsistency percent (): the number of times constraint (3) is violated divided by the number of times the antecedent is true.
@@
 Further, we can use inference in conjunction with this iterative prompting strategy.
-Tables 1 and 2 show the results for QA-SRL and QA-SRL 2.0 respectively.
+Tables 1 and 2 show the results for QA-SRLwiki and QA-SRL 2.0 respectively.
 We see that the constrained models produce near-consistent gains over their unconstrained counterparts.
@@
 Post-inference, the answer to the first question is changed to “lifelong career in medicine”, resulting in the correct structure.
-| System | Head() | Head() | () |
+| System | Headq() | Heads() | () |
 | T5-3B | 63.84 | 36.17 | 34.30 |
@@
 | + constraints | 43.34 | 18.48 | 0 |
-| Flan-T5-XL | 59.47 | 33.60 | 49.79 |
+| Flan-T5-XLitr | 59.47 | 33.60 | 49.79 |
 | + constraints | 61.11 | 36.17 | 0 |
 | GPT-4 | 80.46 | 62.93 | 5.61 |
-| System | Head() | Head() | () |
+| System | Headq() | Heads() | () |
 | T5-3B | 68.57 | 46.35 | 39.06 |
@@
 | + constraints | 66.02 | 48.68 | 0 |
-| Flan-T5-XL | 68.84 | 49.29 | 42.11 |
+| Flan-T5-XLitr | 68.84 | 49.29 | 42.11 |
 | + constraints | 71.97 | 55.41 | 0 |
@@
 For SRL, we compare the small, base, large, 3 billion, and 11 billion T5 models.
-Figure 4 shows the Head metric over T5 model sizes for QA-SRL 2.0, along with the inconsistency percent.
+Figure 4 shows the Heads metric over T5 model sizes for QA-SRL 2.0, along with the inconsistency percent.
 For coreference, we compare the large, 3 billion and 11 billion variants of the Macaw model.
@@
 Interestingly, with QA-SRL 2.0, a constrained model of a certain size often yields comparable, if not better, results as an unconstrained model of the immediately bigger model, showing that constraints might give performance gains equivalent to using a bigger model.
-Appendix C shows model size results for the other datasets.
+Appendix C shows results for the other datasets.
 6.2 Few-shot Prompting
@@
 We consider the latter.
-We conduct few-shot experiments for the SRL task on the QA-SRL (Figure 6) and QA-SRL 2.0 (Figure 7).
-We observe a clear increase in performance with increasing shots for QA-SRL.
+We conduct few-shot experiments for the SRL task on the QA-SRLwiki (Figure 6) and QA-SRL 2.0 (Figure 7).
+We observe a clear increase in performance with increasing shots for QA-SRLwiki.
 Recall that Flan T5, in a zero-shot setting for this dataset, underperforms compared to T5.
-Hence, the gains are not surprising: only the 5-shot constrained system beats the zero-shot constrained T5 model (38.75 Head).
+Hence, the gains are not surprising: only the 5-shot constrained system beats the zero-shot constrained T5 model (38.75 Heads).
 The model also becomes more consistent with more examples.
@@
 Appendix A Detailed SRL Results
-| System | Exact() | Exact() | Head() | Head() | () |
+| System | Exactq() | Exacts() | Headq() | Heads() | () |
 | T5-3B | 38.80 | 12.65 | 63.84 | 36.17 | 34.30 |
@@
 | + constraints | 33.12 | 9.52 | 43.34 | 18.48 | 0 |
-| Flan-T5-XL | 37.57 | 12.32 | 59.47 | 33.60 | 49.79 |
+| Flan-T5-XLitr | 37.57 | 12.32 | 59.47 | 33.60 | 49.79 |
 | + constraints | 43.80 | 17.25 | 61.11 | 36.17 | 0 |
 | GPT-4 | 62.84 | 37.07 | 80.46 | 62.93 | 5.61 |
-| System | Exact() | Exact() | Head() | Head() | () |
+| System | Exactq() | Exacts() | Headq() | Heads() | () |
 | T5-3B | 54.12 | 30.49 | 68.57 | 46.35 | 39.06 |
@@
 | + constraints | 57.06 | 37.23 | 66.02 | 48.68 | 0 |
-| Flan-T5-XL | 55.11 | 32.86 | 68.84 | 49.29 | 42.11 |
+| Flan-T5-XLitr | 55.11 | 32.86 | 68.84 | 49.29 | 42.11 |
 | + constraints | 61.34 | 41.17 | 71.97 | 55.41 | 0 |
@@
 B.1 Zero Shot Coreference Results
-The complete zeros shots results for Macaw-3B and Flan T5-XL models are given in Table 8.
+The complete zero shot results for Macaw-3B and Flan T5-XL models are given in Table 8.
 One might notice that the Flan model is highly inconsistent as compared to the Macaw model.
@@
 Examining the predictions, we saw that the value of the divisor for the Macaw models is roughly four times that of the Flan models.
-This implies that Flan tends to predict ‘No’ more often.
-| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
+This implies that Flan-T5 tends to predict ‘No’ more often.
+| Dataset | System | F1() | MUC() | B3() | CEAFe() | CoNLL() | () |
 | Single | 13.32 | 63.93 | 41.67 | 18.94 | 41.51 | 0 | |
@@
 Furthermore, intermediate context between the sentences can be pivotal in establishing otherwise loosely connected sentences appearing a few sentences apart.
-This begs the question about how much, if any, does providing the entire document help.161616For Ontonotes and GENIA, the document size went beyond the input size which is why we do not test along this dimension.
+This begs the question about how much, if any, does providing the entire document help.161616For Ontonotes and GENIA, the document size went beyond memory which is why we do not test along this dimension.
 On a different note, mentions can be highlighted with asterisks(as used for WSC in Raffel et al.
@@
 We see that results are comparable across all context styles.
-| Dataset | System | F1() | MUC() | B() | CEAF() | CoNLL() | () |
+| Dataset | System | F1() | MUC() | B3() | CEAFe() | CoNLL() | () |
 | Rel | 46.26 | N/A | N/A | N/A | N/A | 48.56 | |
@@
 Appendix C Additional Model Size Experiments
-The analysis over model sizes for QA-SRL is given in Figure 8.
+The analysis over model sizes for QA-SRLwiki is given in Figure 8.
 We see similar analysis as observed with QA-SRL 2.0 dataset.
@@
 In the iterative prompting template, we prompt the model one question at a time such that the answer for every preceding prompt is included in subsequent prompts.
-Assume a predicate in a sentence to have questions: {q,q,…,q}.
-In the first step, we prompt the model for q using the prompt template mentioned in the previous sub-section.
-Next, we follow with the prompt the model for q such that the question and the generated answer for q are included in the prompt.
+Assume a predicate in a sentence to have questions: {q1,q2,…,qn}.
+In the first step, we prompt the model for q1 using the prompt template mentioned in the previous sub-section.
+Next, we follow with the prompt the model for q2 such that the question and the generated answer for q1 are included in the prompt.
 Consequently, we use a different prompt template that contains this pair and an instruction that asks the answers to not overlap.
-Subsequently, the prompt for q will contain the question-answer pair q and q, and so on.
+Subsequently, the prompt for q3 will contain the question-answer pair q1 and q2, and so on.
 An example is shown in Figure 11.
+We use the order provided in the QA-SRL datasets for our experiments.
 D.1.2 GPT-4 Prompt Template
@@
 | Dataset | Gen. step | Inf. Step |
-| QA-SRL | 48.49 | 2.18 |
+| QA-SRLwiki | 48.49 | 2.18 |
 | QA-SRL 2.0 | 945.22 | 32.68 |
@@
 Appendix F SRL Inference Worked-out Example
+| Rank | Role | ||
+| a | b | c | |
+| 1 | “Elrond” (2) | “Aragorn” (5) | “Aragorn the sword” (5) |
+| 2 | “Elrond gave” (1) | “Elrond” (3) | “the sword” (4) |
+| Rank | Path | Path length |
+| 1 | 0 | |
+| 2 | 0 | |
+| 3 | 0 | |
+| 4 | 0 | |
+| 5 | 0 | |
+| 6 | 0 | |
+| 7 | 1 | |
+| 8 | 1 | |
 We will consider the example mentioned in Figure 2 with toy values to illustrate the inference process.
@@
 For evaluation purposes, we treat them as incorrect answers.
-| Rank | Path | Path length |
-| 1 | 0 | |
-| 2 | 0 | |
-| 3 | 0 | |
-| 4 | 0 | |
-| 5 | 0 | |
-| 6 | 0 | |
-| 7 | 1 | |
-| 8 | 1 | |
-| Rank | Role | ||
-| a | b | c | |
-| 1 | “Elrond” (2) | “Aragorn” (5) | “Aragorn the sword” (5) |
-| 2 | “Elrond gave” (1) | “Elrond” (3) | “the sword” (4) |